# FlexiBit: Fully Flexible Precision Bit-parallel Accelerator Architecture for Arbitrary Mixed Precision AI

### v1

@@ 3.3. Primitive Generator
We can observe that the multiplication process (step 3 in Figure 3) consists of (1) computing AND for each operand bit pair and (2) bit shifting and addition. FlexiBit’s PE also follows those steps. For the first step, we design a hardware module named ”Primitive Generator.” We term the individual AND results for each activation-weight bit pair as ”Primitive.” As notation, we refer to the primitive generated between the j-th bit of activation A and i-th bit of weight as $P_{A}(i,j)$, which is defined as $P_{A}(i,j)=A_{i}\cdot W_{j}$ where $\cdot$ refers to the bit-wise AND operator.
As inputs, Primitive Generator receives a set of activation and weight mantissa stored in activation and weight mantissa registers (red registers in Figure 4), respectively. Primitive Generator parses the bits in the activation and weight mantissa registers to generate primitives in the order and layout consumed by FBRT, which will be discussed in Subsection 3.4. For the implementation, we take the same approach as the Separator discussed in Subsection 3.2: utilize small crossbars and keep a reasonable width to constrain the hardware costs while providing performance benefits.  Figure 5 illustrates an example operation of the Primitive Generator. Individual crossbars for weight and activation deliver desired operands to each position, which generates primitives in the ascending order of indices from right to left.
@@
Figure 5. An example of multiplication primitive generation when we have two 3-bit activations and two 3-bit weights. (a) Primitives required for an example multiplication tile in FlexiBit. (b) Primitive generator operation (generation + arrangement) for the example in (a).
@@ 3.4. FBRT: A Flexible-Bit Reduction Tree
Figure 6. The architecture of FBRT and two example operations, inspired by MAERI (Kwon et al., 2018). (a) Multiplication of power-of-2 width activation (2 bits) and weight (4 bits). In power-of-2 cases, additional links (i.e., links between two adjacent nodes in the same level that do not share the same parent node) are not utilized. (b) Multiple multiplications of non-power-of-2 width operands: activation is 3-bit-wide.
@@
FBRT is the key enabler for the flexible-precision bit-parallel multiplication. FBRT approaches the multiplication operation as a spatial shift-add operation and aggregates primitive bits in a tree architecture inspired by inspired by Augmented Reduction Tree (ART) of MAERI (Kwon et al., 2018). FBRT extends the ART to the bit level, which performs concatenation, shifting, and addition of bits with flexible routing of intermediate results. Like ART in MAERI, FBRT includes additional links, which are links between neighboring nodes in the same level that do not share the same parent node. Also, the link bandwidth in the tree gradually increases as the level gets closer to the top (root node). FBRT expects the primitives are sorted in terms of activation indices and then weight indices, as illustrated in Figure 5, which is managed by Primitive Generator. The primitives are fed into the leaf nodes and traverse the pre-configured tree to simultaneously generate multiple multiplication results. The tree nodes have four operating modes with variants within the mode to implement the core functionality, which we discuss next.
Figure 7. Switch modes overview. For $concat\_add$ mode, only one case is shown for simplicity.
@@
Switch Modes. We summarize the switch modes of FBRT in Figure 7. As shown in the table, FBRT’s node switches support four modes with variants depending on the position of the node (left or right to the additional link) and the number of operands a switch handles. The operation of the switches consists of routing, concatenation, shifting, and addition, as shown in the operation description in Figure 7. We will discuss two operation examples to understand how switches enable bit-parallel flexible precision multiplication.
Operation Example.  Figure 6 shows two examples of FBRT operations: (a) one on power-of-two bit-width operands, and (b) the other involving non-power-of-two activations. As shown in Figure 6 (a), if the bitwidth of all operands are power of twos, no additional links are utilized. In that example, primitives are concatenated in the first level and added in the following levels with proper shifting for implementing the floating point multiplication described in Figure 6 Step 3. Unlike this example, additional links are involved if the operand bit widths are non-power-of-twos, as shown in Figure 6 (b). In such a case, additional links play key roles enabling concurrent processing of multiple multiplications. Note that the shifting and addition operations need to be carefully carried out for correct operation, unlike ART of MAERI (Kwon et al., 2018) just focused on the concurrent reduction of irregular-width dot products.
@@
Control Logic. Since the data precision and format are expected to be consistent within each layer, we can utilize a compiler to generate the control signals (i.e., switch modes and shift amounts for additions). We describe the algorithm below in a Python-style pseudo code:
@@
2//oids: An array that stores corresponding output IDs for each bit position in the primitive register.
3//sids: An array that contains corresponding segment IDs for each bit position in the primitive register.
@@
17 if j % 2 == 1: // right case
18 if oids[i-1][j*2]+1 == get_AL_oids[i][j]:
@@
22 else: // left case
23 // Similar logic as right case
24 else: // if SIDs of children SIDs do not match
25 if j % 2 == 1: // right case
26 if oids[i-1][j*2]+1 == get_AL_oids[i][j]:
27 if sids[i-1][j*2]+1 == get_AL_sids[i][j]:
@@
33 else: // left case
34 // Simlar logic as right case
35 else: // if OIDs does not match
36 if j %2 == 1: // right case
37 if isvalid(oids[i-1][j*2]):
38 switch_mode[i][j] = D2
39 else:
40 switch_mode[i][j] = D1
41 else: // left case
42 // Similar logic as left case
Code 3: A pseudo code for FBRT control signal generation
The control signal generation algorithm above simulates the operation in a compiler from the leaf nodes to the top (root node), filling switch modes for each switch in a level from right to left. To determine the switch mode, the algorithm utilizes the operation ID (OID), which refers to the ID of different multiplications, and segment ID, which refers to the ID of rows of intermediate results shown in Step 3 of  Figure 3. Note that a compiler handles the control signal generation, which does not incur runtime overheads.
Handling implicit 1. The FP formats assume implicit 1 in the mantissa for the efficiency of representation. One way to directly support implicit 1 is by adding 1s when we generate primitives. However, this approach leads to a significantly larger number of primitives. For example, if we multiply two mantissa with 2-bit and 3-bit representations, we need 2 $\times$ 3 = 6 primitives. However, if we include implicit 1s, we need (2+1) $\times$ (3+1) = 12 primitives, which is double). To avoid excessive overhead from implicit ones, we utilize implicit 1s’ key characteristics: values are always one. Figure 8 presents our approach to handle implicit 1s. We first generate intermediate outputs without implicit ones using FBRT. After that, in step 1, we exploit the property that left-most values of segments are original weight bits (e.g., W0 and W1 appear in the left-most position of segments from the top in  Figure 8 step 1). That is, we can shift the original weight and add it to the results of FBRT, P_FBRT. Afterwards, we take the same approach for the activation, whose original values are listed in the bottom segment. Note that the left-most bit of the bottom segment is always 1 since it is the primitive that corresponds to the implicit 1s of activation and weight.
@@
3.5. Flexible Bit Exponent Adder
@@ 7 FBEA_control[i] = 0
Code 4: A pseudo code for Flexible Bit Exponent Adder control logic
@@ 3.7. Concat-Shift Tree (CST)
Figure 10 illustrates an example operation of CST with five three-bit mantissas. The difference in the exponent computed by ENU is used to determine the shift amount of each level. The control logic generation follows a similar approach as FBRT, tracking the ID of individual mantissa and concatenating them if values from left and right children belong to the same mantissa ID. If a value from the Additional link belongs to the same mantissa as well, a tree node switch performs three-way concatenation. Note that the shift amount computed from ENU for individual mantissa is applied when performing concat-shift operation.
Figure 10. An example operation of Concat-Shift Tree (CST) on five 3-bit mantissas.
@@ 3.8. Accumulation and Normalization Unit (ANU)
Based on the results from CST, the Accumulation and Normalization Unit (ANU) performs addition of incoming partial outputs, normalizes the exponent based on the results of addition, and shifts the mantissa considering (1) the implicit 1, (2) newly normalized exponent, and (3) target output precision. We adopt the FBEA and ENU to implement the core functionalities of ANU. The control follows the same style as FBEA and ENU.
@@ 4. FlexiBit: Fully-flexible Precision Accelerator
Figure 11 depicts an overview of our fully-flexible precision/format bit-parallel accelerator, FlexiBit. FlexiBit follows the general architecture of AI accelerators at the high level but it contains new architectures for flexible precision and format in the PE and bit-packing and -unpacking unit.
@@
Figure 12. Output Stationary Dataflow of FlexiBit. We map outputs spatially on the PE array. Data distribution on dimension K is done temporally. Depending on the NoC bandwidth, we can send more than one row/column data on dimension K to each PE to increase data reuse.
@@ 4.1. Bit Packing and Unpacking Unit (BPU)
Storing and accessing such non-power-of-two precision data introduces a major challenge to the memory and network-on-chip. One simple approach is to store each value in a fixed-sized block whose size is a power of two by padding zeros to the remaining bits (e.g., store FP6 data into an 8-bit memory block with padding). Although this approach simplifies control logic, it is not ideal for on-chip memory efficiency. Therefore, we implement a crossbar-based bit-packing and -unpacking unit (BPU) as illustrated in Figure 13 to support condensed bit-packed memory layout in the accelerator. The BPU is activated if the host memory utilizes the zero-padded memory layout. BPU is also responsible for propagating metadata to the accelerator controller indicating the starting address, bit precision information, and number of data elements.
Figure 13. The operation example in the BitPacking Unit with FP6.
@@ 4.2. Dataflow
Although FlexiBit requires an outer-product style computation in each PE for efficient flexible precision support, the dataflow in the PE array level is user-configurable.  Figure 12 illustrates two example dataflows supported by FlexiBit adopted for evaluation in the output-stationary (OS) and weight-stationary (WS) style. As long as each tile fits into each PE’s local memory, users can also configure tile sizes. One key enabler for the flexible dataflow is the two-dimensional bus-based network-on-chip architecture illustrated in Figure 12, which allows data transfer to any PE from each SRAM.
@@ 5.1. Evaluation Settings
LLM Workloads. We use three recent language models: Bert-base-uncased (Kenton and Toutanova, 2019), Llama-2-7b (Touvron et al., 2023), and GPT-3 (Brown et al., 2020). We list the hyperparameters of the evaluated models in Table 4.
@@
Hardware Configurations:. We simulate the accelerators in four different configurations: Mobile-A, Mobile-B, Cloud-A, and Cloud-B. We list the details of the configurations in Table 3. We apply the default hardware design parameters listed in Table 1, which was selected after cost-benefit analysis. We deploy weight- and output-staionary dataflow styles illustrated in Figure 12 for our design. We set the design parameters to the default values listed in Table 1, which are optimized after exploring performance and area trade-off.
@@ 5.2. Methodology
We implement our design in RTL using System Verilog and synthesize the design using NanGate 15nm Technology (Martins et al., 2015) with Synopsys Design Compiler. We implement a cycle-accurate simulator designed to model our RTL implementation. The simulator is validated against RTL simulation with 96% accuracy on Bert-base and 99% accuracy on Llama-2-7b, as shown in Figure 14. We utilize a state-of-the-art energy estimation framework, Accelergy (Wu et al., 2019), for evaluating energy, with the synthesis results of our RTL implementation.
@@ 5.3. Evaluation Results
We highlight latency, energy, and performance per area results to discuss the benefits of FlexiBit.
@@
Figure 15. Latency of Bert-Base, Llama-2-7b Llama-2-70b, and GPT-3 models on different act/ weight precisions. Input sequence is 2048. Results are depicted for (a) Mobile-A configuration, (b) Mobile-B configuration, (c)Cloud-A configuration, and (d)Cloud-B configuration accelerator scales.
5.3.1. Latency Comparison

### v2

@@ 3.3. Primitive Generator
We can observe that the multiplication process consists of (1) computing AND for each operand bit pair and (2) bit shifting and addition. FlexiBit’s PE also follows those steps. For the first step, we design a hardware module named ”Primitive Generator.” We term the individual AND results for each activation-weight bit pair as ”Primitive.” As notation, we refer to the primitive generated between the j-th bit of activation A and i-th bit of weight as $P_{A}(i,j)$, which is defined as $P_{A}(i,j)=A_{i}\cdot W_{j}$ where $\cdot$ refers to the bit-wise AND operator.
As inputs, Primitive Generator receives a set of activation and weight mantissa stored in activation and weight mantissa registers (red registers in Figure 2), respectively. Primitive Generator parses the bits in the activation and weight mantissa registers to generate primitives in the order and layout consumed by FBRT, which will be discussed in Subsection 3.4. Since (1) the mantissa bit-widths can be different for both inputs and activations and (2) we want the primitives of each operation to be alongside each other for FBRT, different patterns can occur when parsing the bits. For the implementation, we take the same approach as the Separator discussed in Subsection 3.2: utilize small crossbars and keep a reasonable width to constrain the hardware costs while providing performance benefits.
Figure 3 (c) illustrates an operation example of Primitive Generator for $BW_{M}(A)=3$ and $BW_{M}(W)=2$. To compute the multiplication given in Figure 3 (c), Primitive Generator needs to generate the cross-product of each activation and weight bit. We use two crossbars for weights and activations and match the corresponding operands for each primitive. Note that we generate the primitive in the ascending order of weight/activation bit indices in a packed format, as shown in  Figure 3 (c). Maintaining the order is crucial for the operation of Flexible-Bit Reduction Tree (FBRT), which is the key enabler for the flexible precision/format multiplication. Note that we handle the implicit-1 bit in another module to prevent the exponential enlargement of the crossbar and FBRT size, which will be discussed in Subsection 3.4.
@@ 3.4. FBRT: A Flexible-Bit Reduction Tree
FBRT is the key enabler for the flexible-precision bit-parallel multiplication. FBRT approaches the multiplication operation as a spatial shift-add operation and aggregates primitive bits in a tree architecture inspired by inspired by Augmented Reduction Tree (ART) of MAERI (Kwon et al., 2018). Our contribution beyond the ART is (1) the fine-grained (bit-level) reconfigurability, (2) complex node operations (add, concat, shifting, and their combinations), and (3) control modes that coordinate complex node operations. What we mainly adopt from ART is the fat-tree-like structure and additional links, which are links between neighboring nodes in the same level that do not share the same parent node. FBRT expects the primitives are sorted in terms of activation and weight indices, as illustrated in Figure 3 (d), which is managed by Primitive Generator. The primitives traverse pre-configured tree from leaf nodes to simultaneously generate multiple multiplication results. The tree nodes have four operating modes with variants within the mode to implement the core functionality, which we discuss next.
Figure 4. Microarchitecture switch and possible modes overview. Left and Right nodes refer to the tree node with neighbor link on the left and right side, respectively. The brackets ”{” and ”}” refer to concatenation.
@@
Switch Modes. We summarize the switch modes of FBRT in Figure 4. As shown in the table, FBRT’s node switches support four modes with variants depending on the position of the node (left or right to the additional link) and the number of operands a switch handles. The operation of the switches consists of routing, concatenation, shifting, and addition, as shown in the operation description in Figure 4. To understand how those switches are used to enable bit-parallel flexible precision multiplication, we discuss an operation example next.
Operation Example.  Figure 3 (d) shows an example operation of FBRT on the multiplication between FP6 activations and FP5 weights In the example, red arrows refer to the use of additional links, which is crucial for the bit level flexibility. Various switch mode labels on each tree node show how the complex node operations are combined to enable the flexible bit arithmetic. For example, Out[1] is generated by concatenating two incoming primitives from dependent nodes, left-shifting it by one, and adding it with the other incoming primitive from the additional link ({P(2,0), P(1,0), P(0,0} + ((concat (({ P(2,1),P(1,1)}), P(1,0))) ¡¡ 1). Likewise, primitives are progressively concatenated and added after aligning the bit positions using shifting. Although the control logic on switches handles complex operations, the logic is controlled by a simple finite state machine (FSM) using compiler-generated control signals, which we discuss next.
@@
Control Logic. Since the data precision and format are expected to be consistent within each layer, we can utilize a compiler to statically generate the control signals. We describe the algorithm below in a Python-style pseudo code:
@@
2//oids: array - Output IDs for each bit position in the primitive reg.
3//sids: array - Segment IDs for each bit position in the primitive reg.
@@
17 if oids[i-1][j*2]+1 == get_N_oids[i][j]:
@@
21 else: // if children SIDs do not match
22 if oids[i-1][j*2]+1 == get_N_oids[i][j]:
23 if sids[i-1][j*2]+1 == get_N_sids[i][j]:
@@
29 else: // if children OIDs do not match
30 switch_mode[i][j] = D
Code 3: Pseudo code of FBRT control signal generation
The control signal generation algorithm above simulates the operation in a compiler from the leaf nodes to the top (root node), filling switch modes for each switch in a level from right to left. To determine the switch mode, the algorithm utilizes the operation ID (OID), which refers to the ID of different multiplications, and segment ID, which refers to the ID of rows of intermediate results shown in Figure 5. Note that a compiler handles the control signal generation, which does not incur runtime overheads.
Optimization for the implicit 1. FP formats assume implicit 1 in the mantissa for the efficiency of representation. One way to directly support implicit 1 is by adding 1s when we generate primitives. However, this approach leads to a significantly larger number of primitives. For example, if we multiply two mantissa with 2-bit and 3-bit representations, we need 2 $\times$ 3 = 6 primitives. However, if we include implicit 1s, we need (2+1) $\times$ (3+1) = 12 primitives, which is double). To avoid excessive overhead from implicit ones, we utilize its key characteristics: the value is always one. Figure 5 presents our approach. We first generate intermediate outputs without implicit ones using FBRT. After that, in step 1, we exploit the property that left-most values of segments are original weight bits (e.g., W0 and W1 appear in the left-most position of segments from the top in  Figure 5 step 1). That is, we can shift the original weight and add it to the results of FBRT, P_FBRT. Afterwards, we take the same approach for the activation, whose original values are listed in the bottom segment. Note that the left-most bit of the bottom segment is always 1 since it is the primitive that corresponds to the implicit 1s of activation and weight.
@@
3.5. Flexible Bit Exponent Adder (FBEA)
@@ 7 FBEA_control[i] = 0
Code 4: Pseudo code of FBEA control logic
@@ 3.7. Concat-Shift Tree (CST)
Figure 7 (a) illustrates an example operation of CST with three-bit mantissas. The difference in the exponent computed by ENU is used to determine the shift amount of each level. The control logic generation follows a similar approach as FBRT, tracking the ID of individual mantissa and concatenating them if values from left and right children belong to the same mantissa ID. If a value from the Additional link belongs to the same mantissa as well, a tree node switch performs three-way concatenation. Note that the shift amount computed from ENU for individual mantissa is applied when performing concat-shift operation.
Figure 7. (a) The Concat-Shift Tree (CST) architecture and an example operation on 3-bit mantissas. (b) Node microarchitecture and modes in CST.
@@ 3.8. Accumulation and Normalization Unit (ANU)
Based on the results from CST, the Accumulation and Normalization Unit (ANU) adds incoming partial outputs, normalizes the exponent based on the results of addition, and shifts the mantissa considering (1) the implicit 1, (2) newly normalized exponent, and (3) target output precision. We adopt the FBEA unit and its control mechanism to implement the core functionalities of ANU.
@@ 4. FlexiBit: Fully-flexible Precision Accelerator
Figure 8 depicts an overview of our fully-flexible precision/format bit-parallel accelerator, FlexiBit. FlexiBit follows the general architecture of AI accelerators at the high level but it contains new architectures for flexible precision and format in the PE and bit-packing and -unpacking unit. FlexiBit’s control signals are generated by a compiler since they are deterministic when the precisions are determined. The control signal generation algorithm also can be implemented in a hardware controller, which can be used for dynamic control of the hardware. However, as the granularity of precision/format change is coarse-grained in current practice (layer/operator or model level), we utilize compiler and store the control signals in the control and state registers (CSRs) as showin in  Figure 8. With this approach, the controller’s area overhead is negligible ( 0.2% of the total accelerator area). Note that the control signals can be broadcasted to PEs if the precision changes in the layer or model granularity, which minimizes the reconfiguration cost (less than 100 cycles; note that each GEMM consumes millions-billions of cycles).
@@ 4.1. Bit Packing and Unpacking Unit (BPU)
Storing and accessing such non-power-of-two precision data introduces a major challenge to the memory and network-on-chip. One simple approach is to pad zeros with the remaining bits (e.g., store FP6 data into an 8-bit memory block with padding). Dispite simplified control logic, this approach is not ideal for on-chip memory efficiency. Therefore, we implement a crossbar-based bit-packing unit (BPU) as illustrated in Figure 3 (a), to support condensed bit-packed memory layout in the accelerator. The BPU is activated if the host memory utilizes the zero-padded memory layout. The incoming data is 64 bits with zero-padded data when the offchip interface is 64-bit wide. The controller utilizes precision information (i.e. 6 in this example) and a $start\_idx$ register. Each useful bit in the $i$-th position of the input is mapped to $j$’th position of the output. $j$ is calculated using this formula: $j=start\_idx+i-(\lfloor i/8\rfloor\times(8-precision)$. Figure 3 (a) demonstrates an example with FP6 activation. First six bits on the activation are mapped with the same index. 7th and 8th bits of input are masked and has no effect on the output vector. Input bits of 9 to 14 are mapped to output bits of 7 to 12. In the next data arrival, $start\_idx=start\_idx+precision*8$ is applied and the crossbar maps each input bit using the same method. A double buffering is considered to handle packing overflow properly and transferring the data to the SRAM. BPU is also responsible for propagating metadata to the accelerator controller, indicating the starting address, bit precision information, and number of data elements.
Bit-packing vs Padding. Padded data formats are preferred in system software (OS, runtime, firmware, etc.) due to their data address alignment requirements, which is crucial for the functionality and efficiency of sub-components of system software such as memory allocator. However, padded data involves redundant zeros, which leads to inefficiency. By removing the redundancy, packed data layout reduces the on-chip traffic, memory efficiency, and compute unit utilization. We’ll discuss the overhead in  Subsection 5.3.4.
@@ 4.2. Dataflow
Although FlexiBit PE requires an outer-product style GEMM computation for efficiency, the dataflow in the PE array level is fully flexible in FlexiBit, as long as each compute tile fits into each PE’s local memory. The key enabler for the flexible dataflow is the two-dimensional bus-based NoC, which allows data transfer to any PE from each SRAM. In the evaluation, we select the best dataflow style among weight- and output-stationary dataflows in the PE array level, setting the tile size to be the maximum possible for on-chip memory size.
@@ 5.1. Evaluation Settings
LLM Workloads. We use four language models: Bert-base-uncased (Kenton and Toutanova, 2019), Llama-2-7b (Touvron et al., 2023), Llama-270b (Touvron et al., 2023), and GPT-3 (Brown et al., 2020). We list the hyperparameters of the evaluated models in Table 3.
@@
Hardware Configurations:. We simulate the accelerators in four different configurations: Mobile-A, Mobile-B, Cloud-A, and Cloud-B. We list the details of the configurations in Table 2. We apply the default hardware design parameters listed in Table 1, which was selected after cost-benefit analysis. We deploy weight- and output-staionary dataflow styles for our design. We set the design parameters to the default values listed in Table 1, which are optimized after exploring performance and area trade-offs.
@@ 5.2. Methodology
We implement our design in RTL using System Verilog and conduct Post-PnR analysis using NanGate 15nm Technology (Martins et al., 2015) and Cadence Innovus. We implement a cycle-accurate simulator designed to model our RTL implementation. The simulator is validated against RTL simulation with 96% accuracy on Bert-base and 99% accuracy on Llama-2-7b, as shown in Figure 9. We utilize a state-of-the-art energy estimation framework, Accelergy (Wu et al., 2019), for evaluating energy, with the synthesis results of our RTL implementation. We estimate DRAM read/write energy based on previously reported data (O’Connor et al., 2017).
@@ 5.3. Evaluation Results
Figure 10. Latency of Bert-Base, Llama-2-7b Llama-2-70b, and GPT-3 models on different precisions. Input sequence is 2048. Results are depicted for (a) Mobile-A configuration, (b) Mobile-B configuration, (c)Cloud-A configuration, and (d)Cloud-B configuration accelerator scales.
5.3.1. Latency and Energy Comparison
Our experiments demonstrate that FlexiBit’s performance varies for Output-Stationary (OS) and Weight-Stationary (WS) for different accelerator scales and workload models, which is aligned with previous works on dataflow (Parashar et al., 2019; Kwon et al., 2019). The OS dataflow parallelizes M and N dimensions and reuse partial outputs K times. The WS dataflow parallelizes K and N dimensions and reuse weights M / TileSz(M) times. Their performance and efficiency varies depending on different data reuse counts, as formulated above, and degree of parallelization determined by the interplay between tensor shapes, tile sizes, and accelerator array sizes. In the evaluation, we leverage the dataflow flexibility of FlexiBit and report results based on the best dataflow between WS and OS for each experiment. Note that the dataflow preference can be quickly determined using our performance model.